I could use help on learning whare to place ordered pairs in quadratic equations. I have been ought for a long time because of a illness, so any help is greatly appreciated. my teacher asks : Given the points on the graoh, find the particular equation of the quadratic function containing the given ordered pairs, write the equation in standard form y=ax^2 + bx + c.............
One set of ordered pairs I am given is : (1,6),(3,26),(-2,21).......Thanks for the help, peter

Answer by venugopalramana(3286)

GIVEN ORDERED PAIRS (x,y)
x y
1 6
3 26
-2 21
y=ax^2+bx+c substituting the 3 ordered pairs value for x and y in this eqn.we get
a+b+c=6 I
9a+3b+c=26 II
4a-2b+c=21 III
II-I 8a+2b=20 or 4a+b=10 IV
II-III 5a+5b=5 or a+b=1 V
IV-V 3a=9 or a=3
V 3+b=1 or b=-2
I 3-2+c=6 or c=5
hence y = 3x^2-2x+5